A bullet of mass 0.05kg travelling horizontally at 30ms^-1 embeds itself in wooden block of mass 0.95kg.if the block is suspended by a light vertical string 1m long calculate the maximum inclination of the string at the horizontal


1
Expert's answer
2021-11-14T17:13:51-0500

Explanations & Calculations


  • The finding of the speed of the block + bullet combined system with the help of conservation of linear momentum.

"\qquad\qquad
\begin{aligned}
\small 0.05kg\times30ms^{-1}+0&=\small (0.05kg+0
.95kg)V\\
\small V&=\small 1.5\,ms^{-1}
\end{aligned}"

  • This kinetic energy is expended as the block-bullet system moves up to the topmost position possible.
  • If the inclination set up with the vertical is "\small \theta," then the height to the block-bullet system from the ground is

"\qquad\qquad
\begin{aligned}
\small h&=\small L-L\cos\theta\\
&=\small 1(1-\cos\theta)
\end{aligned}"

  • Finally, by conservation of mechanical energy,

"\qquad\qquad
\begin{aligned}
\small \frac{1}{2}mV^2+0&=\small 0+(m+M)gh\\
\small \frac{1}{2}. 0.05kg.(1.5ms^{-1})^2&=\small (0.05+0.95).9.8.(1-\cos\theta)\\
\small 1-\cos\theta&=\small 0.00576\\
\small \theta&=\small 6.12^0
\end{aligned}"

  • Therefore, the inclination measured w.r.t the horizontal is

"\qquad\qquad\qquad\small 90-6.12 =83.88^0"